OpenStudy (anonymous):

Find the exact value of the radical expression 2 sqrt(7x) + sqrt(7x) - 3 sqrt(7x)- 7

OpenStudy (anonymous):

pleasee

OpenStudy (luigi0210):

\[2 \sqrt{7x}+\sqrt{7x}-3 \sqrt{7x}-7\] This?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

Okay, all you have to do is add like terms.. replace the sqrt(7x) with an x if that makes it easier for you: 2x+x-3x-7 What do you get?

OpenStudy (anonymous):

-7?

OpenStudy (luigi0210):

There you go :)